# Dynamical Phase Transitions Across Slow and Fast Regimes in a Two-Tone Driven Duffing Resonator

### v1

@@ Abstract
The response of nonlinear resonators to multifrequency driving reveals rich dynamics beyond conventional single-tone theory. We study a Duffing resonator under bichromatic excitation and identify a competition between the two drives, governed by their detuning and relative amplitudes. In the slow-beating regime, where the tones are closely spaced, the secondary drive acts as a modulation that induces dynamical phase transitions between coexisting stationary states. We introduce the cycle-averaged amplitude as an order parameter and map the resulting phase diagram as a function of the drive detuning and amplitude ratio, capturing the pronounced asymmetry observed for blue versus red detuning in experiment. We devise a model to link the onset of these transitions to the resonance properties around the nonlinear stationary mode of the system. Our results provide a framework for controlling driven nonlinear systems, enabling state manipulation, and sensing in nanomechanical, optical, and superconducting circuit platforms.
@@ I Introduction
Our understanding of nonlinear driven-dissipative systems is largely built on their response to single-tone driving. Traditional approaches to analyzing these systems rely on approximations that simplify the dynamics, such as the rotating-wave approximation or timescale separation, which neglect fast oscillating terms [33]. While these approximations can be effective for single-frequency systems, they may not accurately capture the dynamics due to multiple tones. Established perturbative methods like Krylov-Bogoliubov [34, 35], Poincaré-Lindstedt [36] and others, while valuable for weakly-nonlinear single-tone systems or systems involving commensurate frequencies, are insufficient to capture dynamics arising due to incommensurate, low-detuned multi-tone drives where subdominant driving responses are usually treated as linear perturbations.
Recently, Ref. [31] attributed the complex trajectories of a two-tone driven Duffing resonator to topological constraints due to the changing phase-space flow governing the dynamics, positing their existence as a robust structural feature of the system’s phase space. It identified a slow regime where the system closely follows the slowly changing vector flow and exhibits orbits that go around both the attractors. However, these orbits disappear in the fast regime, where the system can no longer respond slowly to the drive. Ref. [31] also proposed a modified model that recreates these dynamics in phase space.
@@
In this work, we propose a model based on the theory of linear susceptibilities to identify and delineate different dynamical regions in parameter space. In order to do this, we first revisit the theory of linear and nonlinear response to build up to the modified equations of motion. We then identify an order parameter to determine the onset of dynamical phase transitions in the stationary response. Using this order parameter, we are able to map the regions with dynamical phase transitions with respect to the detuning and relative strength of the secondary tone. Finally, we introduce our model to delineate the parameter space with dynamical phase transitions. Our results make use of a linear theory based approach to gain new insight into the dynamics of a two-tone driven Duffing resonator. They lay the groundwork for the exploration of multi-tone driven nonlinear systems from a new perspective. These results have implications across fields of physics in sensing and metrology, nonlinear optomechanics, optimized control of qubits, circuit quantum electrodynamics, Floquet-engineering in ultra-cold atoms [29, 37, 38, 21, 39]. Crucially, they are also important beyond physics in the study of tipping points in the context of early warning signs in climate dynamics, ecological systems, and socioeconomic models [4, 40, 41, 42].
@@
The remainder of this paper is structured as follows: In Sec. II, we introduce our model system: a Duffing resonator subject to a bichromatic drive. To build intuition, we first revisit the linear regime of the system in Sec. III, establishing a rotating frame formalism that reveals how the two tones interact in the absence of nonlinearity. In Sec. IV, we review the canonical single-tone driven Duffing resonator, detailing the harmonic balance method (HBM) used for its analysis. In Sec. V, we analyze the slow-beating regime, treating the secondary tone as a slow modulation that can induce dynamical phase transitions, and we introduce an order parameter to map out the resulting phase diagram. In Sec. VI, we develop a model based on a linear-response analogy, refined with nonlinear corrections, to analytically approximate the observed transition boundaries. Finally, we conclude in Sec. VII with a summary of our key findings and an outlook on future research directions.
@@ III Two-tone linear response theory
with $\chi_{\raisebox{-0.90417pt}{$\scriptscriptstyle\Omega_{0}$}}(\omega)=\left(\Omega_{0}^{2}-\omega^{2}-i\Gamma\omega\right)^{-1}$ the susceptibility of the damped harmonic resonator [44], which characterizes the resonator’s inherent frequency-dependent response, see Fig. 1(b). In linear systems, the equation of motion precludes frequency mixing. Consequently, the total response (2) is a superposition of individual responses. The response amplitude (and thus the power) depends on the drive strength $F_{i}$, the resonator susceptibility $\chi_{\raisebox{-0.90417pt}{$\scriptscriptstyle\Omega_{0}$}}(\omega)$, and the detuning; see Fig. 1(b). The response diminishes as the drive frequency deviates from $\Omega_{0}$. Hence, the combined effects of its strength and detuning determine which drive elicits the stronger response.
@@
Figure 1: (a) Potential energy landscape $V(x)$ of a Duffing resonator [cf. Eq. (1)] with positive, vanishing, and negative nonlinearity $\alpha$ (blue, grey, red lines, respectively). Inset: Sketch of a Duffing resonator subject to two different drives (wavering inbound arrows) and dissipation (dashed outbound arrow). (b) Stationary amplitude $x(\omega)$ of a linear resonator in response to two driving tones (vertical solid and dashed lines) [cf. Eq. (2)]. These amplitudes reflect the scaled susceptibility of the linear damped resonator, $\tilde{x}(\omega)=F_{i}\chi_{\raisebox{-0.90417pt}{$\scriptscriptstyle\Omega_{0}$}}(\omega)$ (solid and dashed Lorentzian peaks). Despite larger drive strength in the solid case, the drive in the dashed case yields a higher response, as it is closer to the resonator’s natural frequency. Insets: the rotating frame linear vector flow towards the stationary response [cf. Eq. (5)]. (c) The response of the Duffing resonator to a single drive [cf. Eq. (9)], showing its characteristic bistability regime involving a high and a low amplitude response (pink and blue solid lines, respectively), separated by a saddle (yellow dashed line). We use $\Omega_{0}=2\pi$, $\Gamma=6.2\times 10^{-4}$, $F_{1}=4.81\times 10^{-4}$. (d) The vector flow (solid grey lines) given by Eq. (9) for the driving tone stationed in the bistable region at $\Delta_{10}=-9.02\times 10^{-4}$.
@@
with effective force $\vec{F}=F_{1}[1+h\cos(\Delta_{21}t),-h\sin(\Delta_{21}t)]^{T}$ written in the frame with detuning $\tilde{\Delta}_{01}=\sqrt{\Omega_{0}^{2}-\Omega_{1}^{2}}$. These harmonic equations for the quadratures can be used to visualize the vector flow in a $2D$ rotating phase space, as is measured in experiments with a lock-in amplifier [45]. Closed-loop stationary orbits in the phase space of the laboratory frame appear as stable states, i.e., fixed-point attractors, in the rotating phase space spanned by the quadratures $X(t)$ and $Y(t)$. Additionally, a red-detuned drive results in an anticlockwise vector flow around the attractor, whereas a blue-detuned drive leads to a clockwise vector flow, see insets in Fig. 1(b), lending a notion of chirality to the attractor in the rotating frame [46, 47, 48, 49].
The dynamics of the quadratures, governed by Eq. (5), directly determine the time-dependent squared amplitude of the resonator’s response in the rotating frame, $A^{2}_{\raisebox{-0.90417pt}{$\scriptscriptstyle\Omega_{0}$}}(t,h)$, which quantifies the power measured with a lock-in amplifier at $\Omega_{1}$. To find $A^{2}_{\raisebox{-0.90417pt}{$\scriptscriptstyle\Omega_{0}$}}(t,h)$, we first solve Eq. (5) for $\tilde{X}(\omega)$ and $\tilde{Y}(\omega)$ via Fourier transformation. A further inverse Fourier transform yields (cf. Appendix A for details):
@@ IV Nonlinear single-tone response
While the linear response to two tones already exhibits complex temporal dynamics [cf. Eq. (6)], the introduction of the nonlinearity fundamentally alters the system’s behaviour. For a Duffing resonator driven by a single drive $F_{2}=0$ in Eq. (1)], the nonlinear cubic term induces frequency mixing, i.e., the generation of new frequency components such as harmonics of the input frequency [36]. This precludes an exact analytical solution via Fourier transformation. However, approximate analytical solutions can be obtained via perturbative methods such as the Krylov-Bogoliubov method [35, 34], the Poincaré-Lindstedt method [36], secular perturbation theory [50], or the HBM [51]. These methods rely on separating the timescale dominating the stationary system response from the faster timescales determined by the resonance frequency, or the drive acting on the system.
@@
$A^{2}\left(\frac{\Gamma^{2}}{4}+\left(\frac{3\alpha A^{2}}{4\Omega_{1}}+\frac{\Omega_{1}^{2}-\Omega_{0}^{2}}{\Omega_{1}}\right)^{2}\right)=\frac{F_{1}^{2}}{4\Omega_{1}^{2}}\,.$ (10)
@@
$|F_{lb/hb}|^{2}\!=\!\frac{8}{81\alpha}\!\Big[\!\left(\tilde{\Delta}_{01}^{4}-3\Gamma^{2}\Omega_{1}^{2}\right)^{\frac{3}{2}}\!\\ \pm\tilde{\Delta}_{01}^{2}\left(\tilde{\Delta}_{01}^{4}+9\Gamma^{2}\Omega_{1}^{2}\right)\!\Big]\,$, (11)
and establishes the relationship between the critical drive strength $|F_{lb/hb}|^{2}$ for the low/high amplitude branch to bifurcate, and the bifurcation frequencies $\Omega_{1}$, where the instability occurs. Thus, we can explicitly define the drive strength required to induce a bifurcation at a given frequency, or conversely, the frequencies at which bifurcations occur for a fixed drive strength. Crucially, for fixed ($\tilde{\Delta}_{01},\Gamma,\Omega_{1},\alpha$) these thresholds are not equal: to jump up (low amplitude $\to$ high amplitude) the drive amplitude, $F_{1}$, must exceed the $F_{lb}$, while to jump down (high $\to$ low) it must be reduced below $F_{hb}$. Thus, sweeping the drive amplitude traces a hysteresis loop with direction-dependent switching powers.
@@ V Two-tone nonlinear response
Figure 2: (a) The modulated effective drive amplitude, $F_{\mathrm{eff}}$, over two cycles of the relative phase $\varphi$ for $h=0.36>h_{lb}$. The horizontal lines indicate the bifurcation thresholds, $F_{lb}$ (green) and $F_{hb}$ (orange)[cf. Eq. (11)], and the primary tone’s amplitude $F_{1}$. (b) The stationary amplitude response as a function of drive frequency. The black (grey) curve shows the response at the maximum(minimum), $F_{\mathrm{eff}}^{\mathrm{max(min)}}$ for the modulation shown in (a). The coloured lines show the response to a single drive, i.e for $h=0$ as in Fig. 1(c). The vertical grey line marks the frequency of the primary tone, $\Omega_{1}$. (c) Bifurcation diagram of the response versus the effective drive strength along the drive-modulation cycle [cf. Eq. (10)] for a fixed detuning $\Delta_{10}=-9.02\times 10^{-4}$. The pink(blue) solid line shows the high(low) branch and the yellow dashed line shows the saddle. The orange circle(green square) marks $F_{hb(lb)}$. The vertical grey line marks $|F_{\mathrm{eff}}|=F_{1}$. The arrows show the jump to the higher branch and back as $F_{\mathrm{eff}}$ changes. (d) Quasistatic phase diagram obtained by solving Eq. (12) as a function of the relative drive strength $h$ and phase $\varphi(T)$. The grey bistability region, where both branches exist, is flanked by the low-branch ($h_{lb}$) and high-branch ($h_{hb}$) bifurcation thresholds (solid and dashed black arcs, respectively). The dashed white circle at $h=0.36$ indicates the trajectory for the exemplary vector flows at the black dots in the three distinct regions, shown in panel (e). (e) Exemplary vector flows for $h=0.36$ at three points along the cycle: $\varphi=0.1\pi$ (in the blue region where only the lower branch is stable), $\varphi=0.9\pi$ (in the pink region where only the higher branch is stable), and $\varphi=1.4\pi$ (in the grey bistable region) from bottom to top, respectively. The solid vertical lines show the evolution of the blue(pink) low(high) branch as $\varphi$ varies. The parameters used are $\Omega_{0}=2\pi$, $\Gamma=6.2\times 10^{-4}$, and $F_{1}=4.81\times 10^{-4}$.
@@
Hereafter, we assume the system is initialized in the low-amplitude stationary state of the single-tone problem (i.e., when $F_{2}=0$). The modulated amplitude of $F_{\mathrm{eff}}$ can trigger a jump if $|F_{\mathrm{eff}}|$ is large enough for the system to cross the lower branch bifurcation at $\Omega_{1}$, see Fig. 2(c). Thus, for the jumps to occur, $F_{\mathrm{eff},+}$ as $\varphi$ is varied must satisfy or overshoot the discriminant condition for bifurcation, [cf. Eq. (11)]. This leads to a critical value $h_{lb}$ for the lower branch when $F_{\mathrm{eff},+}=F_{lb}\equiv F_{1}(1+h_{lb})$ exactly satisfies Eq. (11). For $h>h_{lb}$, the system jumps when the low-amplitude attractor loses stability as $\varphi$ is varied and moves to the high-amplitude attractor, see Fig. 2(c). Similarly, the high amplitude branch bifurcates when $F_{\mathrm{eff},-}\leq F_{hb}\equiv F_{1}(1-h_{hb})$ and the system moves back to the low-amplitude attractor.
@@
These inter-attractor jumps can be tracked by looking at the change in the stability of the attractors for varying $\varphi$ and $h$ in Fig. 2(d): for a circular trajectory at fixed $h>h_{lb}$, as $\varphi$ varies and $F_{\mathrm{eff}}$ changes, the system crosses the bifurcation point of the lower branch and consequently jumps to the available stationary branch and moves across stability regions where the low, high or both branches are stable as in Fig. 2(e). This agrees with the recent experiment and analysis presented in Ref. [31]. However, when the beating between the tones becomes faster, the quasistatic picture breaks down, necessitating a more comprehensive analysis to capture the system’s dynamics.
@@
Figure 3: The stationary states calculated from the quasistatic Eq. (12) - pink (blue) solid for high (low) amplitude state, yellow dashed line for the saddle and the corresponding simulation (black line) with $\Delta_{21}=0.42\Gamma$ for (a) $h=0.1$, (b) $h=0.33$ and (c) $h=0.65$.
@@
Building on Ref. [31], the decisive timescale for applying the quasi-static picture is the resonator ringdown, $\tau=1/\Gamma$, which low-pass filters the beat-note modulation at $|\Delta_{21}|$. When $|\Delta_{21}|\ll\Gamma$ (beat period $\gg\tau$), the modulation is quasistatic: the vector-flow evolves slowly and the state continuously follows the stationary branches until the branch bifurcates [Fig. 2(b)]. This enforces completed transitions between the low- and high-amplitude Duffing attractors once per cycle. Conversely, when $|\Delta_{21}|\lesssim\Gamma$ (beat period $\lesssim\tau$), the modulation is fast: dissipation cannot relax the response to the drifting attractor before it merges with the saddle, so the trajectory lags the changing flow and remains trapped near the initial basin, producing small loops without completed inter-attractor jumps despite an unchanged instantaneous vector-flow. Thus, $\tau$ seems to set the boundary between slow transitions in Fig. 2(c) and fast, lagging dynamics. In the following, we show that this initial intuition from Ref. [31] does not depict the full picture.

### v2

@@ Abstract
The response of nonlinear resonators to multifrequency driving reveals rich dynamics beyond conventional single-tone theory. We study a Duffing resonator under bichromatic excitation and identify a competition between the two drives, governed by their detuning and relative amplitudes. In the slow-beating regime, where the tones are closely spaced, the secondary tone acts as a modulation that induces dynamical phase transitions between coexisting stationary states. We introduce the cycle-averaged amplitude as an order parameter and map the resulting phase diagram as a function of the drive detuning and amplitude ratio, capturing the pronounced asymmetry observed for blue versus red detuning between the drives in experiment. We link the onset of these transitions to the resonance properties around the nonlinear stationary mode of the system. We apply a two-tone ansatz to this system, for the first time, to provide a comprehensive map of stationary states. Our results provide a framework for controlling driven nonlinear systems, enabling state manipulation, and sensing in nanomechanical, optical, and superconducting circuit platforms.
@@ I Introduction
Our understanding of nonlinear driven-dissipative systems is largely built on their response to single-tone driving. Traditional approaches to analyzing these systems rely on approximations that simplify the dynamics, such as the rotating-wave approximation or timescale separation, which neglect fast oscillating terms [33]. While these approximations can be effective for single-frequency systems, they may not accurately capture the dynamics due to multiple tones. As such, established perturbative methods like Krylov-Bogoliubov [34, 35], Poincaré-Lindstedt [36] and others, while valuable for weakly-nonlinear single-tone systems or systems involving commensurate frequencies, are insufficient to capture dynamics arising due to incommensurate, low-detuned multi-tone drives where subdominant driving responses are usually treated as linear perturbations.
Recently, the complex trajectories of a two-tone driven Duffing resonator were examined through the lens of the changing phase-space flow [31]. A slow regime was identified where the system follows the flow to encircle coexisting attractors, in contrast to a fast regime where such global orbits are suppressed. While this perspective provides geometric intuition, it relies on a phenomenological reconstruction of the dynamics. A predictive framework that maps the boundaries of these dynamical phase transitions based on the resonator’s response function is currently lacking.
@@
In this work, we develop an effective nonlinear response theory around driven stationary states to predict the thresholds for dynamical phase transitions. To this end, we first revisit the theory of linear and nonlinear response to establish the modified equations of motion. We then introduce an order parameter to map the dynamical phase diagram as a function of the detuning and relative strength of the secondary tone. Crucially, we extract an analytical threshold for the onset of these transitions, providing a vital tool for experimentalists to identify stable operating regimes. Furthermore, we apply a two-tone ansatz for the first time to this system, elucidating the complex dynamical responses beyond the reach of standard approximations. Our results leverage a linear-theory-based approach to gain new insight into the dynamics of a two-tone driven Duffing resonator. These results have implications across fields of physics in sensing and metrology, nonlinear optomechanics, optimized control of qubits, circuit quantum electrodynamics, and Floquet-engineering in ultra-cold atoms [29, 37, 38, 21, 39]. They are also important beyond physics in the study of tipping points in the context of early warning signs in climate dynamics, ecological systems, and socioeconomic models [4, 40, 41, 42].
@@
The remainder of this paper is structured as follows: In Sec. II, we introduce our model system: a Duffing resonator subject to a bichromatic drive. To build intuition, we first revisit the linear regime of the system in Sec. III, establishing a rotating frame formalism that reveals how the two tones interact in the absence of nonlinearity. In Sec. IV, we review the canonical single-tone driven Duffing resonator, detailing the harmonic balance method (HBM) used for its analysis. In Sec. V, we analyze the slow-beating regime, treating the secondary tone as a slow modulation that can induce dynamical phase transitions, and we introduce an order parameter to map out the resulting phase diagram. In Sec. VI, we refine a linear-response-based approach with nonlinear corrections to analytically approximate the observed transition boundaries. In Sec. VII, we apply an extended (two-tone) ansatz to this system for the first time to provide a comprehensive map of stationary states and connect the results with the analytical boundaries in Sec. VI. Finally, we conclude in Sec. VIII with a summary of our key findings and an outlook on future research directions.
@@ III Two-tone linear response theory
with $\chi_{\raisebox{-0.90417pt}{$\scriptscriptstyle\Omega_{0}$}}(\omega)=\left(\Omega_{0}^{2}-\omega^{2}-i\Gamma\omega\right)^{-1}$ the susceptibility of the damped harmonic resonator [44], which characterizes the resonator’s inherent frequency-dependent response, see Fig. 1(b). In linear systems, the equation of motion precludes frequency mixing. Consequently, the total response (2) is a superposition of individual responses. The response amplitude (and thus the power) depends on the drive strength $F_{i}$, the resonator susceptibility $\chi_{\raisebox{-0.90417pt}{$\scriptscriptstyle\Omega_{0}$}}(\omega)$, and the detuning; see Fig. 1(b). The response diminishes as the drive frequency deviates from $\Omega_{0}$. Hence, the combined effects of its strength and detuning determine which drive leads to a stronger response.
@@
Figure 1: The Duffing resonator’s linear response to two drives and nonlinear response to a single-drive. (a) Potential energy landscape $V(x)$ of a Duffing resonator [cf. Eq. (1)] with positive (blue), vanishing (grey), and negative (red) nonlinearity $\alpha$. Inset: Sketch of a Duffing resonator subject to two different drives (wavy inbound arrows) and dissipation (dashed outbound arrow). (b) Stationary amplitude $\tilde{x}(\omega)$ of a linear resonator driven by two tones [cf. Eq. (2)]. The solid and dashed curves represent the scaled susceptibility ${F_{i}}\chi_{\raisebox{-0.90417pt}{$\scriptscriptstyle\Omega_{0}$}}(\omega)$ for different drive strengths $F_{i}$. Note that a weaker drive closer to resonance (dashed) can yield a higher response than a stronger detuned drive (solid). Insets: Rotating frame vector flow [cf. Eq. (5)]; a positively detuned drive induces clockwise motion, while a negatively detuned one induces anti-clockwise motion. (c) Single-tone nonlinear response [cf. Eq. (9)] exhibiting bistability between low (blue) and high (pink) amplitude branches, separated by an unstable saddle (yellow dashed). We use $\Omega_{0}=2\pi$, $\Gamma=6.2\times 10^{-4}$, $F_{1}=4.81\times 10^{-4}$. (d) Phase space vector flow (grey lines) for a drive fixed in the bistable region ($\Delta_{10}=-9.02\times 10^{-4}$), showing two counter-rotating attractors (pink and blue dots) separated by a saddle (yellow cross).
@@
with effective force $\vec{F}=F_{1}[1+h\cos(\Delta_{21}t),-h\sin(\Delta_{21}t)]^{T}$ written in the $\Omega_{1}$-frame, with $\tilde{\Delta}_{01}=\sqrt{\Omega_{0}^{2}-\Omega_{1}^{2}}$. These harmonic equations for the quadratures can be used to visualize the vector flow in a $2D$ rotating phase space, as is measured in experiments with a lock-in amplifier [45]. Closed-loop stationary orbits in the phase space of the laboratory frame appear as stable states, i.e., fixed-point attractors, in the rotating phase space spanned by the quadratures $X(t)$ and $Y(t)$. Additionally, a negatively detuned drive results in an anticlockwise vector flow around the attractor, whereas a positively detuned drive leads to a clockwise vector flow, see insets in Fig. 1(b), lending a notion of chirality to the attractor in the rotating frame [46, 47, 48, 49].
The dynamics of the quadratures, governed by Eq. (5), directly determine the time-dependent squared amplitude of the resonator’s response in the rotating frame, $A^{2}_{\raisebox{-0.90417pt}{$\scriptscriptstyle\Omega_{0}$}}(t,h)=X^{2}(t,h)+Y^{2}(t,h)$, which quantifies the power measured with a lock-in amplifier at $\Omega_{1}$. To find $A^{2}_{\raisebox{-0.90417pt}{$\scriptscriptstyle\Omega_{0}$}}(t,h)$, we first solve Eq. (5) for $\tilde{X}(\omega)$ and $\tilde{Y}(\omega)$ via Fourier transformation. A further inverse Fourier transform yields (cf. Appendix A for details):
@@ IV Nonlinear single-tone response
While the linear response to two tones already exhibits complex temporal dynamics [cf. Eq. (6)], the introduction of the nonlinearity fundamentally alters the system’s behaviour. For a Duffing resonator driven by a single drive $F_{2}=0$ in Eq. (1)], the nonlinear cubic term induces frequency mixing, i.e., the generation of new frequency components such as harmonics of the input frequency [36]. This precludes an exact analytical solution via Fourier transformation. However, approximate analytical solutions can be obtained via perturbative methods such as the Krylov-Bogoliubov method [35, 34], the Poincaré-Lindstedt method [50], secular perturbation theory [2, 51], or the HBM [52]. These methods rely on separating the timescale dominating the stationary system response from the faster timescales determined by the resonance frequency, or the drive acting on the system.
@@
$A^{2}\left({\Gamma^{2}}+\left(\frac{3\alpha A^{2}}{4\Omega_{1}}+\frac{\Omega_{1}^{2}-\Omega_{0}^{2}}{\Omega_{1}}\right)^{2}\right)=\frac{F_{1}^{2}}{\Omega_{1}^{2}}\,.$ (10)
@@
$|F_{\mathrm{lb/hb}}|^{2}\!=\!\frac{8}{81\alpha}\!\Big[\!\left(\tilde{\Delta}_{01}^{4}-3\Gamma^{2}\Omega_{1}^{2}\right)^{\frac{3}{2}}\!\\ \pm\tilde{\Delta}_{01}^{2}\left(\tilde{\Delta}_{01}^{4}+9\Gamma^{2}\Omega_{1}^{2}\right)\!\Big]\,$, (11)
and establishes the relationship between the critical drive strength $|F_{\mathrm{lb/hb}}|^{2}$ for the low/high amplitude branch to bifurcate, and the bifurcation frequencies $\Omega_{1}$, where the instability occurs. Thus, we can explicitly define the drive strength required to induce a bifurcation at a given frequency, or conversely, the frequencies at which bifurcations occur for a fixed drive strength. Crucially, for fixed ($\Delta_{01},\Gamma,\Omega_{1},\alpha$) these thresholds are not equal: to jump up (low amplitude $\to$ high amplitude) the drive amplitude, $F_{1}$, must exceed the $F_{\mathrm{lb}}$, while to jump down (high $\to$ low) it must be reduced below $F_{\mathrm{hb}}$. Thus, sweeping the drive amplitude traces a hysteresis loop with direction-dependent switching.
@@ V Two-tone nonlinear response
Figure 2: The slow modulation picture. (a) The modulated effective drive amplitude, $F_{\mathrm{eff}}$, over two cycles of the relative phase $\varphi$ for $h=0.36>h_{\mathrm{lb}}$. The horizontal lines indicate the bifurcation thresholds, $F_{\mathrm{lb}}$ (green) and $F_{\mathrm{hb}}$ (orange) [cf. Eq. (11)], and the primary tone’s amplitude $F_{1}$ (black). (b) The stationary amplitude response as a function of drive frequency. The black (grey) curve shows the response at the maximum (minimum), $F_{\mathrm{eff}}^{\mathrm{max(min)}}$ for the modulation shown in (a). The coloured lines show the response to a single drive, i.e for $h=0$ as in Fig. 1(c). The vertical grey line marks the frequency of the primary tone, $\Omega_{1}$. (c) Bifurcation diagram against effective drive strength. As $|F_{\mathrm{eff}}|$ oscillates, the system follows the stable branches; arrows indicate the forced jumps between the low (blue) and high (pink) branches when thresholds are crossed. The orange circle (green square) marks $F_{hb\,(lb)}$. The vertical grey line marks $|F_{\mathrm{eff}}|=F_{1}$. (d) Quasistatic phase diagram [cf. Eq. (12)] as a function of the relative drive strength $h$ and phase $\varphi(T)$. The grey region marks bistability, separated by bifurcation thresholds (solid and dashed black arcs) to high- and low-branch only regions. Dashed white circle: trajectory for $h=0.36$. (e) Snapshot of vector flows at the three (black) points along the white trajectory in (d): monostable low (bottom), bistable (middle), and monostable high (top). The solid vertical arcs show the evolution of the blue (pink) low (high) branch as $\varphi$ varies. Parameters: $\Omega_{0}=2\pi$, $\Gamma=6.2\times 10^{-4}$, and $F_{1}=4.81\times 10^{-4}$.
@@
Hereafter, we assume the system is initialized in the low-amplitude stationary state of the single-tone problem (i.e., when $F_{2}=0$). The modulated amplitude of $F_{\mathrm{eff}}$ can trigger a jump if $|F_{\mathrm{eff}}|$ is large enough for the system to cross the lower branch bifurcation at $\Omega_{1}$, see Fig. 2(c). Thus, for the jumps to occur as $\varphi$ is varied, $F_{\mathrm{eff},+}$ must satisfy or overshoot the discriminant condition for bifurcation, [cf. Eq. (11)]. This leads to a critical value $h_{\mathrm{lb}}$ for the lower branch when $F_{\mathrm{eff},+}=F_{\mathrm{lb}}\equiv F_{1}(1+h_{\mathrm{lb}})$ exactly satisfies Eq. (11). For $h>h_{\mathrm{lb}}$, the system jumps when the low-amplitude attractor loses stability as $\varphi$ is varied and moves to the high-amplitude attractor, see Fig. 2(c). Similarly, the high amplitude branch bifurcates when $F_{\mathrm{eff},-}\leq F_{\mathrm{hb}}\equiv F_{1}(1-h_{\mathrm{hb}})$ and the system moves back to the low-amplitude attractor.
@@
These inter-attractor jumps can be tracked by looking at the change in the stability of the attractors for varying $\varphi$ and $h$ in Fig. 2(d): for a circular trajectory at fixed $h>h_{\mathrm{lb}}$, as $\varphi$ varies and $F_{\mathrm{eff}}$ changes, the system crosses the bifurcation point of the lower branch and consequently jumps to the available stationary branch and moves across stability regions where the low, high or both branches are stable as in Fig. 2(e). This agrees with the recent experiment and analysis presented in Ref. [31]. However, when the beating between the tones becomes faster ($|\Delta_{21}|\geq\Gamma$), the quasistatic picture breaks down, necessitating a more comprehensive analysis to capture the system’s dynamics.
@@
Figure 3: The time-domain response $A(T)$ (black curve), simulated over three modulation cycles, attempting to follow the stationary states calculated from the quasistatic Eq. (12) - pink (blue) solid for high (low) amplitude state, yellow dashed line for the saddle, with $\Delta_{21}=0.42\Gamma$ for (a) $h=0.1$, (b) $h=0.33$ and (c) $h=0.65$.
@@
Building on Ref. [31], the decisive timescale for applying the quasi-static picture is the resonator ringdown, $\tau=1/\Gamma$, which low-pass filters the beat-note modulation at $|\Delta_{21}|$. When $|\Delta_{21}|\ll\Gamma$ (beat period $\gg\tau$), the modulation is quasistatic: the vector-flow evolves slowly and the state continuously follows the stationary branches until the branch bifurcates [Fig. 2(b)]. This enforces completed transitions between the low- and high-amplitude Duffing attractors once per cycle. Conversely, when $|\Delta_{21}|\lesssim\Gamma$ (beat period $\lesssim\tau$), the modulation is fast: dissipation cannot relax the response to the changing attractor before it merges with the saddle, so the trajectory lags the changing flow and remains trapped near the initial basin, producing small loops without completed inter-attractor jumps despite an unchanged instantaneous vector-flow with respect to the quasistatic case. Thus, naively, $\tau$ seems to set the boundary between slow transitions in Fig. 2(c) and fast, lagging dynamics. In the following, we show that this initial intuition from Ref. [31] does not depict the full picture.